# An Improved Lower Bound on the Image of the 2-adic Character Map for the Heisenberg Algebra via Modular Linear Differential Equations

### v1

@@ 3.3. Explicit Formulations
Given our description of each $\Delta(q)^{n}\lambda_{n,m}$ as a solution of the third-degree MLDE given in Proposition 3.5, we follow Section 4.3 of [FM16b] in order to describe these in terms of hypergeometric series: In the notation of this paper, we can find that
$r_{1}=2n,\hskip 20.0ptr_{2}=\frac{a}{2}\cdot 2^{m}+\frac{n}{2},\hskip 20.0ptr_{3}=\frac{a}{2}\cdot 2^{m}+\frac{n}{2}+\frac{1}{2}$
@@
$\eta(q)^{8a\cdot 2^{m}+24n}K^{-\frac{a}{3}\cdot 2^{m}+n}\hskip 1.0pt_{3}F_{2}\left(-\tfrac{a}{3}\cdot 2^{m}+n,-\tfrac{a}{3}\cdot 2^{m}+n+\tfrac{1}{3},-\tfrac{a}{3}\cdot 2^{m}+n+\tfrac{2}{3}\right.\\ \left.;-\tfrac{a}{2}\cdot 2^{m}+\tfrac{3}{2}\cdot n+1,-\tfrac{a}{2}\cdot 2^{m}+\tfrac{3}{2}\cdot n+\tfrac{1}{2};K\right),$
where $K=1728/j$. Thus $\Delta(q)^{n}\lambda_{n,m}$ is equal to the above expression up to some constant $C_{n,m,a}$. We assume that $3n<a\cdot 2^{m}$ throughout, which we can do since we will eventually be taking the limit as $m\to\infty$. The hypergeometric series above is a specialization of the form studied in the main theorem of [KT01]: Setting $k=4a\cdot 2^{m}-12n$ and $\lambda=2$ in their notation and ensuring the necessary conditions on these parameters hold, the above expression then gives
(3.9) $\Delta(q)^{n}\lambda_{n,m}=C_{n,m}\Delta(q)^{2n}\left(\sum_{j=0}^{\lfloor(a/3)\cdot 2^{m}-n\rfloor}\frac{(3n-a\cdot 2^{m})_{3j}}{\psi_{j}}E_{4}(q)^{a\cdot 2^{m}-3n-3j}\Delta(q)^{j}\right)$
@@
$\psi_{j}$ $=2^{-12j}(j!)\prod_{l=1}^{j}\left(8l-4a\cdot 2^{m}+12n\right)\left(8l-4a\cdot 2^{m}+12n-4\right)$
(3.10) $=2^{-8j}(j!)(3n-a\cdot 2^{m}+1)_{2j}$
@@
Let $n,m,a\geq 1$. We have
$\lambda_{n,m}=\sum_{j=0}^{\lfloor(a/3)\cdot 2^{m}-n\rfloor}\psi_{j}^{-1}\left(3n-a\cdot 2^{m}\right)_{3j}(15)^{a\cdot 2^{m}}E_{4}(q)^{a\cdot 2^{m}-3n-3j}\Delta(q)^{n+j}.$
@@
(3.11) $\lambda_{n,m}=C_{n,m}\Delta(q)^{n}\left(\sum_{j=0}^{\lfloor(a/3)\cdot 2^{m}-n\rfloor}\frac{(3n-a\cdot 2^{m})_{3j}}{\psi_{j}}E_{4}(q)^{a\cdot 2^{m}-3n-3j}\Delta(q)^{j}\right).$
Note that this shows the $\lambda_{n,m}$ are indeed modular of level one (see Remark 3.2). It remains to show that $C_{n,m}=(15)^{a\cdot 2^{m}}$. Recall from (3.2) that
(3.12) $\lambda_{n,m}=\lambda^{n}\left(E_{4}(q)-2^{4}E_{4}(q^{2})\right)^{a\cdot 2^{m}}+2^{1-12n}\left.\left(\lambda^{-n}\left(E_{4}(q^{2})-E_{4}(q)\right)^{a\cdot 2^{m}}\right)\right\rvert\hskip 2.0ptU_{2}.$
We equate the above two displays and extract the coefficient of $q^{n}$. Set $j=0$ in (3.11) to see that this coefficient is $C_{n,m}$, since $\Delta(q)^{n}=q^{n}+O(q^{n+1})$. In (3.12) only the first term provides a coefficient of $q^{n}$ since $\lambda^{n}=q^{n}+O(q^{n+1})$ also. Thus $C_{n,m}=(1-2^{4})^{a\cdot 2^{m}}=(15)^{a\cdot 2^{m}}$ as desired. ∎
Let us set $a=3$ from now on. Expanding $\Delta(q)=8000G_{4}(q)^{3}-147G_{6}(q)^{2}$ and using the fact that $G_{4}(q)=240E_{4}(q)$, define
@@ 4.2. Convergence of Pre-Images
We will need some results on the $2$-adic properties of the coefficients $c_{n,m,i}$.

### v2

@@ 3.3. Explicit Formulations
Given our description of each $\Delta(q)^{n}\lambda_{n,m}$ as a solution of the third-degree MLDE given in Proposition 3.4, we follow Section 4.3 of [FM16] in order to describe these in terms of hypergeometric series: In the notation of this paper, we can find that
$r_{1}=2n,\hskip 20.0ptr_{2}=\frac{3}{2}\cdot 2^{m}+\frac{n}{2},\hskip 20.0ptr_{3}=\frac{3}{2}\cdot 2^{m}+\frac{n}{2}+\frac{1}{2}$
@@
$\Delta(q)^{2^{m}+n}K^{-2^{m}+n}\hskip 1.0pt_{3}F_{2}\left(-2^{m}+n,-2^{m}+n+\tfrac{1}{3},-2^{m}+n+\tfrac{2}{3}\right.\\ \left.;-\tfrac{3}{2}\cdot 2^{m}+\tfrac{3}{2}\cdot n+1,-\tfrac{3}{2}\cdot 2^{m}+\tfrac{3}{2}\cdot n+\tfrac{1}{2};K\right),$
where $K=1728/j$. Thus $\Delta(q)^{n}\lambda_{n,m}$ is equal to the above expression up to some constant $C_{n,m}$. We assume that $n<\cdot 2^{m}$ throughout, which we can do since we will eventually be taking the limit as $m\to\infty$. The hypergeometric series above is a specialization of the form studied in the main theorem of [KT01]: Setting $k=12\cdot 2^{m}-12n$ and $\lambda=2$ in their notation and ensuring the necessary conditions on these parameters hold, the above expression then gives
(3.9) $\Delta(q)^{n}\lambda_{n,m}=C_{n,m}\Delta(q)^{2n}\left(\sum_{j=0}^{2^{m}-n}\frac{(3n-3\cdot 2^{m})_{3j}}{\psi_{j}}E_{4}(q)^{3\cdot 2^{m}-3n-3j}\Delta(q)^{j}\right)$
@@
$\psi_{j}$ $=2^{-12j}(j!)\prod_{l=1}^{j}\left(8l-12\cdot 2^{m}+12n\right)\left(8l-12\cdot 2^{m}+12n-4\right)$
(3.10) $=2^{-8j}(j!)(3n-3\cdot 2^{m}+1)_{2j}$
@@
Let $n,m\geq 1$. We have
$\lambda_{n,m}=\sum_{j=0}^{2^{m}-n}\psi_{j}^{-1}\left(3n-3\cdot 2^{m}\right)_{3j}15^{3\cdot 2^{m}}E_{4}(q)^{3\cdot 2^{m}-3n-3j}\Delta(q)^{n+j}.$
@@
(3.11) $\lambda_{n,m}=C_{n,m}\Delta(q)^{n}\left(\sum_{j=0}^{2^{m}-n}\frac{(3n-3\cdot 2^{m})_{3j}}{\psi_{j}}E_{4}(q)^{3\cdot 2^{m}-3n-3j}\Delta(q)^{j}\right).$
Note that this shows the $\lambda_{n,m}$ are indeed modular of level one. It remains to show that $C_{n,m}=15^{3\cdot 2^{m}}$. Recall from (3.2) that
(3.12) $\lambda_{n,m}=\lambda^{n}\left(E_{4}(q)-2^{4}E_{4}(q^{2})\right)^{3\cdot 2^{m}}+2^{1-12n}\left.\left(\lambda^{-n}\left(E_{4}(q^{2})-E_{4}(q)\right)^{3\cdot 2^{m}}\right)\right\rvert\hskip 2.0ptU_{2}.$
We equate the above two displays and extract the coefficient of $q^{n}$. Set $j=0$ in (3.11) to see that this coefficient is $C_{n,m}$, since $\Delta(q)^{n}=q^{n}+O(q^{n+1})$. In (3.12) only the first term provides a coefficient of $q^{n}$ since $\lambda^{n}=q^{n}+O(q^{n+1})$ also. Thus $C_{n,m}=(1-2^{4})^{3\cdot 2^{m}}=15^{3\cdot 2^{m}}$ as desired. ∎
Expanding $\Delta(q)=8000G_{4}(q)^{3}-147G_{6}(q)^{2}$ and using the fact that $G_{4}(q)=240E_{4}(q)$, define
@@ 4.2. Convergence of Pre-Images
The purpose of this subsection is to develop results necessary for the proof of Proposition 4.6, namely that for all $n,m\geq 1$, the pre-images $v_{n,m}$ satisfy $\nu_{2}(v_{n,m})\geq-6n$. These results involve a close examination of the $2$-adic properties of the coefficients $c_{n,m,i}$. In contrast to [BF24, FM25, FM23], we avoid obtaining explicit expressions for the $v_{n,m}$ in the round-bracket formalism. In fact, in Lemma 4.3, we prove a useful tool for assessing the valuation of the pre-images $\alpha_{r}\beta_{s}$ (cf. (4.1)) for any $r,s\geq 2$. This will be used repeatedly towards the proof of Proposition 4.6.